Installing and Setting up Virtual Box

1 Introduction

For your dfs lab, we want you to have a standardized environment in which to test your code without hogging the use of the school computers. As such we’ve customized an image for you to use with the virtualizer of your choice. Howerver, we strongly recommend that you use VirtualBox as it will minimize any differences among virtualizers that you may encounter. We will not provide support for other virtualizers, so use them at your own risk!

2 Downloading and Installation

Visit:

http://www.virtualbox.org/wiki/Downloads

and download the version that corresponds to the system that you have. Go through the installation (just agree to everything) and start up VirtualBox.

3 Setting it up

First, make sure you have the image that we’ve made available to you (440.vdi).

To create your virtual machine:

1. Go to File → Virtual Media Manager

2. In the manager, click “Add”

3. Locate the image we gave you and select it.

4. It should appear in the list of hard disks.

5. Click OK to exit the Virtual Media Manager

Then:

1. Click “New”

2. Click “Next”

3. For “Name”, enter a name of your choosing. “440” will work just fine.

4. Under Operating System, choose “Linux”.

5. Under Version, choose “Ubuntu”.

6. Select the amount of base memory you want the virtual machine to have. The amount the wizard recommends is fine.

7. Choose “Use existing hard disk” and make sure 440.vdi is selected.

8. Click Finish.

To start up your virtual machine, simply click on it and hit Start. The username for the machine is 440User and the password is “systems”. You do not have to log in when using the virtual machine.